I do not understand how they got their answer to this problem:
Write an equation of the line containing the given point and parallel to the given line
(7,-7) 6x-5y=3
My answer was y=6/5x-35/5
The correct answer is y=6/5x-77/5 How do they get 77 out of this? I am so lost..
Thank you

Answer by Alan3354(69443)

Write an equation of the line containing the given point and parallel to the given line
(7,-7) 6x-5y=3
My answer was y=6/5x-35/5
The correct answer is y=6/5x-77/5 How do they get 77 out of this?
-----------------------
m = 6/5
Use y = mx + b and the point to find b
-----
-7 = (6/5)*7 + b = 42/5 + b
b = -77/5
--> y = (6/5)x - 77/5